OpenStudy (anonymous):

The volume of a rectangular box is given by the formula v=LWH. The volume of the box is 120Cu. units. find the width and length. ( L=x+2, width=x, ht=5

OpenStudy (anonymous):

Ooh, I personally think that this is a fun one :)

OpenStudy (anonymous):

Okay, so your equation is v=lwh

OpenStudy (anonymous):

You know that the volume is 120, the length is (x+2), the height is 5 so 120=(x+2)(x)(5)

OpenStudy (anonymous):

sorry, my internet disconnected

OpenStudy (anonymous):

So start by multiplying. x*5=5x 120=(x+2)(5x)

OpenStudy (anonymous):

You can distribute the 5x through the (x+2) 5x(x+2) 5x^2 +10x 120=5x^2 +10x

OpenStudy (anonymous):

Can you see how this is turning into a quadratic?

OpenStudy (anonymous):

Like a quadratic, you need it to be equal to zero, so move the 120 over. 0=5x^2 +10x -120

OpenStudy (anonymous):

Wait- do you know what I mean by "quadratic?"

OpenStudy (anonymous):

Dede- if you leave me hanging, I'm leaving :P

OpenStudy (anonymous):

I'm watching how you coming about the process.

OpenStudy (anonymous):

you mean ax2 + bx + c = 0

OpenStudy (anonymous):

gandalfwiz p/s come back!